OpenStudy (anonymous):

Two boats traveling the same direction leave a harbor at noon. After 2 hr, they are 40 miles apart. If one boat travels twice as fast as the other, find the rate of each boat

OpenStudy (anonymous):

20 mi/hr and 40 mi/hr

OpenStudy (anonymous):

\[2y-2x=40\]\[y=2x\]Solve the equations

OpenStudy (anonymous):

Distance = (rate)(time)

OpenStudy (anonymous):

y = 10

OpenStudy (anonymous):

xy=-10

OpenStudy (anonymous):

x=1/2y

OpenStudy (anonymous):

no... u can check my answer by analyzing the problem. After traveling for 2 hrs, they are 40 mi apart. \[40 X 2 = 80\]\[20 x 2 = 40. \]\[80-40=40\]\[Then 40 is twice of 20.\]

OpenStudy (anonymous):

for the slowest car 20 yes I get it now

OpenStudy (anonymous):

boat 2(r-40) =4(r+40)